Grade 12Physical Chemistry

pressure at depth 10ft of swimming pool will be what if the atm pressure is 1 atm.?

rutuja vibhute
10 Years agoGrade 12

2 Answers

Vikas TU
10 Years ago
 
pressure at depth 10ft of swimming pool will be what if the atm pressure is 1 atm
will be:
P = pgh + 1
   = 1*10*3.048 + 1
   = 31.48 atm
 
 
naveen
10 Years ago
 
pressure at depth 10ft of swimming pool will be what if the atm pressure is 1 atm
will be:
P = pgh + 1
   = 1*10*3.048 + 1
   = 31.48 atm